OpenStudy (anonymous):

For a driveway that is 32 feet wide, 12 yards long, and 9 inches thick, how many tons of Portland cement concrete is required if the unit weight of PCC is assumed to be 150pcf?

OpenStudy (anonymous):

32*12*9 = 3456 so there's 3,456 cubed feet of cement needed. I'm assuming that is 150 pounds per 150 cubed feet so 3456 * 150 = 518400 there are 2000 pounds in a ton so... 518400/2000 = 259.2 you would need 259.2 tons of cement to fill the driveway